OpenStudy (anonymous):

what is the x-intercept and y-intercept for this equation: y=-1/2x+3, please help and explain why, thank you

OpenStudy (anonymous):

\[y=-1/2x+3\]

OpenStudy (anonymous):

to find x int plug in 0 for y to find y int plug in 0 for x

OpenStudy (anonymous):

y int is y= -1/3 x int dne

OpenStudy (anonymous):

wait there is an x int x = 0 is just a asymptote

OpenStudy (anonymous):

This is currently in slope-intercept form. y=mx+b, where m is the slope and b is the y intercept. So first off, your y-intercept would be 3. An x intercept would be where y=0, so we can do that now. 0=-1/2x+3 subtract 3 from both sides. -3=-1/2x Multiply by the reciprocal of -1/2, which is -2/1. 6=x. So the y intercept is (0,3) and the x intercept is (6,0)

OpenStudy (anonymous):

okay thank you! that help a lot.

OpenStudy (anonymous):

oops, helped

OpenStudy (anonymous):

actually, it wouldn't be (6,0) it would be (-6,0)

OpenStudy (anonymous):

Yes, it is 6,0. 0=(-1/2)x+3 What multiplied by -1/2 would equal 0? 6. \[-\frac{ 1 }{ 2 }x+3\] \[-\frac{ 1 }{ 2 }(6)+3\] \[0=-3+3\]